# Missing Metals in DQ Stars; a Compelling Clue to their Origin††thanks: Dedicated to the memory of Prof. G. Fontaine (1948–2019)

### v2

@@ Abstract
White dwarf stars whose spectra exhibit molecular carbon features have been known for over 60 years, when the first pair were discovered by carefully constructed spectrophotometry (Greenstein & Matthews, 1957; Bell, 1962). Owing to instrumentation limits on wavelength at that time, only the weaker C (1,0) feature near 4670 Å was observable (Eggen & Greenstein, 1965; Greenstein, 1970), and not the stronger C (0,0) feature with bandhead near 5165 Å that would later be found in many more stars (e.g. Oke 1974; Liebert 1977; Wesemael et al. 1993). It has been understood for nearly half a century that white dwarfs exhibiting C Swan bands (or other carbon features) in their spectra – denoted as spectral class DQ (McCook & Sion, 1987) – have atmospheres that are dominated by helium, with only trace abundances of carbon, and little or no hydrogen (Bues, 1973; Grenfell, 1974).
@@
Fontaine et al. (1984) first mentioned the striking problem that is the subject of this work, “… it is interesting to speculate on the reasons why metals and carbon still appear to be mutually exclusive in the atmospheres of cool helium-rich white dwarfs.” The DQ stars are around 9–10 per cent of the local population as defined by the 20 pc sample, which is well constrained if not complete (Hollands et al., 2018), and thus are not outliers or unusual by any means. Indeed, more than 60 years after the first discovery, this fundamental question about their nature remains unanswered, despite recent and ongoing work with far superior data and modeling, and based on hundreds of DQ and DZ stars (Koester & Kepler, 2019; Coutu et al., 2019; Koester et al., 2020). Recent work has shown that it is possible to suppress or mask the Swan bands in stars with sufficient metal abundances resulting from external pollution, but not in all cases (Blouin, 2022), and hence this conundrum remains.
@@ 2 Spectroscopy and metal abundances
Notes and references: (1) Blouin & Dufour (2019); (2) Weidemann & Koester (1995); (3) Sayres et al. (2012); (4) Coutu et al. (2019); (5) Giammichele et al. (2012); (6) Dufour et al. (2005); (7) Subasavage et al. (2017); (8) Farihi et al. (2013a); (9) C band detected in the optical for the first time; (10) carbon detected only in ultraviolet.
@@ 2.1 Target selection and observations
The observed DQ stars are listed in Table 1, and taken from the literature based on favorable brightness and position in the sky. Science targets were selected to have K, and thus are primarily classical DQ stars with C Swan bands, where trace atmospheric carbon is attributed to the dredge up of core material via successful modeling in helium-rich white dwarfs (Pelletier et al., 1986; Koester et al., 2020). These cooler DQ white dwarfs were selected because they represent the bulk of all objects in this spectral class, and form a modest but significant fraction of the local white dwarf population (Hollands et al., 2018). One known DQZ star (0913+103; Wegner et al. 1985; Kleinman et al. 2013) was included in the study in order to have at least one confident detection of Ca ii H and K, and to search for other metal species and thereby better assess the rare, externally polluted DQ stars.
@@ 3.1 Accretion rates: UVES
In stark contrast to their shared effective temperatures and luminosities, in terms of metal accretion rates, the DQ and DZ stars show no overlap whatsoever. It may appear that there are no DZ stars with abundances and inferred accretion rates as low as the DQ(Z) objects, but there has not yet been any search for weak metal lines in DC white dwarfs using large telescopes and high-resolution spectroscopy. The conventional detection of DZ stars is straightforward using low-resolution spectroscopy of modest quality, based on the fact that cool helium atmospheres are largely transparent owing to a lack of major opacity sources (Koester, 2009). Notably, the Ca ii lines in vMa 2 were detected with photographic plate technology over 100 years ago (van Maanen, 1917), and, as can be seen in Figure 1, using an 8 m class telescope and a high-resolution spectrograph is clearly overkill, even for [Ca/He] ( g s). For context, DZ stars with metal abundances 30 times smaller than vMa 2 are known, and the corresponding Ca ii line strengths are readily detected using the resolving power of SDSS spectroscopy (Dufour et al., 2007; Coutu et al., 2019).
The use of the VLT and UVES, together with a sample size of 23 targets, demonstrates conclusively that the pollution in DQZ stars never approaches the levels typically attained in DAZ, DBZ, and DZ stars. For systems that are likely in accretion-diffusion equilibrium (i.e. a steady state), instantaneous accretion rates can exceed g s (Farihi et al., 2016), and roughly speaking, that is 5600 times higher than the rates inferred for polluted DQ white dwarfs. And while steady-state accretion cannot be assumed for helium-rich atmospheres, if the time-averaged mass accretion rates are compared between the DQZ and the DZ stars, the top accretors in each class are separated by a factor of 18 000(!). Furthermore, most DQ stars have only upper limit metal abundances and accretion rates, and therefore the gulf between these two populations is vast.
@@ 3.2 Pollution frequency: SDSS
This example demonstrates that, owing to the intrinsic width of the metal absorption features in cool, helium atmosphere white dwarfs – exemplified by vMa 2 in Figure 1 – the detection of DQZ stars using SDSS spectroscopy is broadly comparable to their detection with UVES, at least to the same degree as found in 0913+103. Using this DQZ white dwarf as a benchmark, it can be inferred that abundances and time-averaged accretion rates as low as [Ca/He] and g s are likely detectable, and thus absent in the DQ stars in the SDSS. This conclusion adds considerable weight to the stark difference in pollution between the DQ and DZ white dwarfs.
@@
As can be seen from the Table, there is a general coherence in the picture of polluted white dwarfs among the spectral classes DA(Z), DB(Z), and DC/Z, where this is now a well-understood story of dynamically active planetary systems in the post-main sequence (Veras, 2016). Most of the aforementioned pollution frequencies and accretion rate inferences belong to studies more than a decade old, and in the intervening years those interpretations have been strengthened. After cooling for 20 Myr or so (during which the sensitivity to external pollution is poor, owing to hotter atmospheres; Koester et al. 2014; Barstow et al. 2014), a given white dwarf may exhibit metals at the observed frequency as a fraction of its remaining cooling lifetime (i.e. a duty cycle), or a subset of the entire population may stay constantly polluted, while others are never enriched by remnant planetary systems.
@@ 4.1 Lack of core dredge-up in DZ white dwarfs?
That being said, for a range metal abundances in DZ white dwarfs it has been demonstrated that atmospheric carbon bands, that might otherwise result from dredge-up, can be suppressed or masked (Blouin, 2022). Updated modeling shows that the C Swan band opacity decreases significantly with increasing metal abundance associated with external pollution, and thus, in principle, a star that appears spectrally of type DQ, may in principle transform into spectral type DZ. While certainly plausible for a range of parameters, roughly one quarter of the modeled DZQ spectra appear detectable with sufficient S/N. Such an empirical test should discover if this potential mechanism actually operates.
@@
It is well known that some DQ white dwarfs exhibit carbon features in the optical, while others do so only in the ultraviolet, and that this can be ascribed to differences in helium mass fractions rather than effective temperature (Weidemann & Koester, 1995; Blouin et al., 2023). In fact, as noted in Table 1, several of the DQ targets were found to have an optical C feature for the first time in their UVES spectra (0806–661, 1708–147, 1837–619), while others only exhibit carbon features in the ultraviolet (1917–077, 2147+280, 2317–173) despite the new, highly-sensitive observations. Therefore, it is possible or even likely that there is a continuous distribution of helium layer masses (Koester et al., 2020), where thinner layers will lead to optical carbon features, modestly thicker layers result in ultraviolet carbon features only, and if sufficiently thick, no carbon will be effectively dredged.
@@ 4.2 Lack of circumstellar material orbiting DQ stars
The dearth of detected circumstellar material orbiting DQ white dwarfs may appear – superficially – redundant with the lack of atmospheric pollution, as there is a strong connection established between these two hallmarks of evolved planetary systems. As a rule for white dwarfs, photospheric heavy elements require ongoing or recent accretion; that is, pollution implies closely-orbiting circumstellar material, where it is understood that the bulk of debris disks are simply undetectable owing to sensitivity limits (Rocchetto et al., 2015; Bonsor et al., 2017). But the converse is not necessarily true, and planetary matter can exist in a wide range of orbits and sizes that would not necessarily lead to the chemical enrichment of the host star. Little is known about the architecture of planetary systems orbiting polluted white dwarfs outside of a few tens of R (Xu et al., 2013; Farihi et al., 2014), but there are at least two examples of the aforementioned scenario: there is a giant planet candidate, transiting the non-polluted white dwarf WD 1856+534 (spectral type DC; Vanderburg et al. 2020), and a 2500 AU, planetary-mass companion orbiting WD 0806-661 (spectral type DQ, Table 1; Luhman et al. 2012).
@@ 4.3 Observed lack of close DQ+dM binaries
Because DC and DQ white dwarfs share similar stellar properties – with the likely exception of the atmospheric helium mass fraction – one would expect a comparable number of each in the SDSS catalog, adjusted for their relative occurrence rates as field stars. Within the 20 pc sample, there are 29 DC and 15 DQ white dwarfs (Hollands et al., 2018), and if their frequency in unresolved (including short-period) binaries were similar, one would expect close to 35 DQ spectral types within the composite binary catalog. Therefore, the lack of DQ stars in this catalog is striking. While identifying a weak C 5165 Å feature may be difficult in low S/N spectra, and in the presence of flux from an M dwarf companion, many DQ white dwarfs exhibit deep and strong Swan bands that are unmistakable (Dufour et al., 2005).
@@
To demonstrate the plausibility of this atmospheric transformation, consider a K, pure helium atmosphere white dwarf, which will manifest as a DB spectral type, and where the fully mixed outer layer has mass M(Koester, 2009; Koester et al., 2020). If this white dwarf accretes solar composition wind from a companion at a modest rate of M yr (Pyrzas et al., 2012; Parsons et al., 2012; Ribeiro et al., 2013), hydrogen will come to dominate by number within a fraction of a Myr. With sufficient sensitivity, a hydrogen(-enriched) atmosphere white dwarf accreting stellar wind should appear as a DAZ spectral type, and there are several well-documented cases among post-common envelope binaries (Zuckerman et al., 2003; Debes, 2006; Tappert et al., 2011).
However, for the aforementioned, pure helium atmosphere case, once the convection zone deepens due to further cooling, any superficial hydrogen layer may be easily overwhelmed. By the time such a white dwarf has cooled to K, the size of the fully mixed layer should be M, where this applies both to DQ white dwarfs and their helium atmosphere DC star counterparts (Koester, 2009; Koester et al., 2020). With convection zones of this depth, it would take several Gyr or longer for the atmosphere to become dominated by hydrogen that is captured from a companion stellar wind, with a strong dependence on the binary separation (i.e. wind capture rate). Therefore, it is expected that some DC, DZ, and DQ white dwarfs should be hosts to post-common envelope, main-sequence companions. Indeed, there are a handful of such systems known to host DC stars (Nebot Gómez-Morán et al., 2011; Parsons et al., 2013), and at least one candidate DZ accreting metals in a stellar wind from an M dwarf (Fajardo-Acosta et al., 2016). But there are none with a DQ white dwarf.
@@ 4.4 Apparently low DQ white dwarf masses
The left-hand panel of Figure 5 plots the mass distribution for the sample of 221 classical DQ white dwarfs discussed in Section 3.2 (Koester & Kepler, 2019). The masses of DQ stars have been studied recently by two independent teams, based on sources with reliable parallax data from Gaia, finding that their average mass is roughly 0.05 M lower than those found for DA and DB stars (Coutu et al., 2019; Bédard et al., 2022), and also lower than DZ and DC white dwarfs (Koester & Kepler, 2019). There have been suggestions that there may be model imperfections that prevent better agreement between DQ stars and other white dwarfs, but to date, no other studies have considered enhanced mass loss that may result from binary evolution.
It should be noted that two iconic DQ white dwarfs in the 20 pc sample have masses derived independently of models, and these values are also lower than the typical M found for spectral type DA and DB white dwarfs (Genest-Beaulieu & Bergeron, 2019). In fact, the dynamical mass of Procyon B (0.59 M) is in modest tension with the overall age estimated for this well-studied binary system (Liebert et al., 2013; Bond et al., 2015). Adopting a comprehensive, recent, and empirical initial-to-final mass relation, the predicted mass of Procyon B should be closer to 0.66 M (Cummings et al., 2018). Similarly, for the planet-hosting binary GJ 86, the DQ white dwarf has a mass of 0.54 M (Zeng et al., 2021). And while it is likely this system is relatively old (either thick disk or somewhat younger; Fuhrmann et al. 2014), any recent initial-to-final mass relation will yield a progenitor that should not have evolved off the main sequence in a Hubble time. Therefore, it is possible that the masses of DQ stars are actually low compared to other white dwarfs, and not a result of model shortcomings, but as a consequence of their evolutionary history.
Among the 221 DQ white dwarfs plotted in Figure 5, there are 23 objects with derived masses M, where 16 of these have , and thus excellent astrometric data. It is often the case that (apparently) over-luminous white dwarfs are two unresolved stars with comparable brightness, and therefore Gaia DR3 was searched in a 5 arcsec radius around the positions of these 23 candidate low-mass DQ stars; none were found to be resolved into two sources. However, two were found to have renormalized unit weight error (RUWE) suggestive of a possible or likely non-single star: SDSS J094058.34+384350.2 (RUWE ) and SDSS J125245.44+194311.2 (RUWE ). Thus, the inferred low masses of these 23 DQ white dwarfs may be partly accounted for as binary light sources, but the bulk currently lack indications of duplicity.
@@
Given that all DQ stars must have atmospheric carbon by definition, and thus carbon cores, it may seem oxymoronic to speculate on the possibility of low-mass members of this class. Owing to the finite age of the Galaxy, white dwarfs with M cannot yet have formed by single star evolution. Indeed the majority of low-mass white dwarfs are found to be single- or double-lined binaries (Marsh et al., 1995; Brown et al., 2011; Rebassa-Mansergas et al., 2011), whose evolution is thought to be terminated prior to helium ignition, and thus such stars would have helium cores as opposed to carbon cores.
However, it is possible to form a low-mass carbon core in stars with M, where helium ignites non-degenerately (Bauer & Kupfer, 2021). Binary interactions would then be required to truncate stellar evolution and further core growth, in order to later be observed as a low-mass white dwarf with a carbon core. While speculative, this is a distinct possibility for some DQ white dwarfs, as their binary properties are only weakly constrained at present, and their molecular bands are relatively insensitive to radial velocity changes induced by companions, especially compared to white dwarfs with atomic spectral lines. The confirmation of any DQ white dwarf with M would be a strong indication of a binary origin.
The expectations for single star evolution are illustrated in the right-hand panel of Figure 5, where main-sequence progenitor masses have been calculated using the previously mentioned initial-to-final mass relation. Given the low masses of the DQ white dwarfs, it is thus not surprising that the predicted main-sequence predecessors are also of relatively low mass, but this important detail has so far escaped attention in previous studies (Coutu et al., 2019; Koester & Kepler, 2019; Bédard et al., 2022). The age of thin disk stars should be less than around 8 Gyr (Bensby et al., 2014; Kilic et al., 2017; Sharma et al., 2019), but Figure 5 indicates that most single star parents of DQ white dwarfs would have main-sequence lifetimes longer than this. For the DQ stars plotted in the Figure, K and M, where white dwarf cooling would add an additional 1.2 Gyr to their inferred total ages. This picture would imply that DQ stars are dominated by thick disk and halo members, which, as shown below, is inconsistent with their kinematical and companion metallicity properties.

### v3

@@ Abstract
White dwarf stars whose spectra exhibit molecular carbon features have been known for over 60 years, when the first pair were discovered by carefully constructed spectrophotometry (Greenstein & Matthews, 1957; Bell, 1962). Owing to instrumentation limits on wavelength at that time, only the weaker C2 (1,0) feature near 4670 Å was observable (Eggen & Greenstein, 1965; Greenstein, 1970), and not the stronger C2 (0,0) feature with bandhead near 5165 Å that would later be found in many more stars (e.g. Oke 1974; Liebert 1977; Wesemael et al. 1993). It has been understood for half a century that white dwarfs exhibiting C2 Swan bands (or other carbon features) in their spectra – denoted as spectral class DQ (McCook & Sion, 1987) – have atmospheres that are dominated by helium, with only trace abundances of carbon, and little or no hydrogen (Bues, 1973; Grenfell, 1974).
@@
Fontaine et al. (1984) first mentioned the striking problem that is the subject of this work, “… it is interesting to speculate on the reasons why metals and carbon still appear to be mutually exclusive in the atmospheres of cool helium-rich white dwarfs.” The DQ stars are around 9–10 per cent of the local population as defined by the 20 pc sample, which is well constrained if not complete (Hollands et al., 2018), and thus are not outliers or unusual by any means. Indeed, more than 60 years after the first discovery, this fundamental question about their nature remains unanswered, despite recent and ongoing work with far superior data and modeling, and based on hundreds of DQ and DZ stars (Koester & Kepler, 2019; Coutu et al., 2019; Koester et al., 2020). In response to an earlier version of the study presented here (Farihi et al. 2022; arXiv:2208.05990v1), it has been shown that it is possible to suppress or mask the Swan bands in stars with sufficient metal abundances resulting from external pollution, but not in all cases (Blouin, 2022), and hence this conundrum remains.
@@ 2 Spectroscopy and metal abundances
Notes and references: (1) Blouin & Dufour (2019); (2) Weidemann & Koester (1995); (3) Sayres et al. (2012); (4) Coutu et al. (2019); (5) Giammichele et al. (2012); (6) Dufour et al. (2005); (7) Subasavage et al. (2017); (8) Farihi et al. (2013a); (9) C2 band detected in the optical for the first time; (10) carbon detected only in ultraviolet.
@@ 2.1 Target selection and observations
The observed DQ stars are listed in Table 1, and taken from the literature based on favorable brightness and position in the sky. Science targets were selected to have K, and thus are primarily classical DQ stars with C2 Swan bands, where trace atmospheric carbon is attributed to the dredge up of core material via successful modeling in helium-rich white dwarfs (Pelletier et al., 1986; Koester et al., 2020). These cooler DQ white dwarfs were selected because they represent the bulk of all objects in this spectral class, and form a modest but significant fraction of the local white dwarf population (Hollands et al., 2018). One known DQZ star (0913+103; Wegner et al. 1985; Kleinman et al. 2013) was included in the study in order to have at least one confident detection of Ca ii H and K, and to search for other metal species and thereby better assess the rare, externally polluted DQ stars.
@@ 3.1 Accretion rates: UVES
In stark contrast to their shared effective temperatures and luminosities, in terms of metal accretion rates, the DQ and DZ stars show no overlap whatsoever. It may appear that there are no DZ stars with abundances and inferred accretion rates as low as the DQ(Z) objects, but there has not yet been any search for weak metal lines in DC white dwarfs using large telescopes and high-resolution spectroscopy. The conventional detection of DZ stars is straightforward using low-resolution spectroscopy of modest quality, based on the fact that cool helium atmospheres are largely transparent owing to a lack of major opacity sources (Koester, 2009). Notably, the Ca ii lines in vMa 2 were detected with photographic plate technology over 100 years ago (van Maanen, 1917), and, as can be seen in Figure 1, using an 8 m class telescope and a high-resolution spectrograph is clearly overkill, even for [Ca/He] ( g s-1). For context, DZ stars with metal abundances 30 times smaller than vMa 2 are known, and the corresponding Ca ii line strengths are readily detected using the resolving power of SDSS spectroscopy (Dufour et al., 2007; Coutu et al., 2019).
The use of the VLT and UVES, together with a sample size of 23 targets, demonstrates conclusively that the pollution in DQZ stars never approaches the levels typically attained in DAZ, DBZ, and DZ stars. For systems that are likely in accretion-diffusion equilibrium (i.e. a steady state), instantaneous accretion rates can exceed g s-1 (Farihi et al., 2016), and roughly speaking, that is 5600 times higher than the rates inferred for polluted DQ white dwarfs. And while steady-state accretion cannot be assumed for helium-rich atmospheres, if the time-averaged mass accretion rates are compared between the DQZ and the DZ stars, the top accretors in each class are separated by a factor of 18 000(!). Furthermore, most DQ stars have only upper limit metal abundances and accretion rates, and therefore the gulf between these two populations is vast.
@@ 3.2 Pollution frequency: SDSS
This example demonstrates that, owing to the intrinsic width of the metal absorption features in cool, helium atmosphere white dwarfs – exemplified by vMa 2 in Figure 1 – the detection of DQZ stars using SDSS spectroscopy is broadly comparable to their detection with UVES, at least to the same degree as found in 0913+103. Using this DQZ white dwarf as a benchmark, it can be inferred that abundances and time-averaged accretion rates as low as [Ca/He] and g s-1 are likely detectable, and thus absent in the DQ stars in the SDSS. This conclusion adds considerable weight to the stark difference in pollution between the DQ and DZ white dwarfs.
@@
As can be seen from Table 2, there is a general coherence in the picture of polluted white dwarfs among the spectral classes DA(Z), DB(Z), and DC/Z, where this is now a well-understood story of dynamically active planetary systems in the post-main sequence (Veras, 2016). Most of the aforementioned pollution frequencies and accretion rate inferences belong to studies more than a decade old, and in the intervening years those interpretations have been strengthened. After cooling for 20 Myr or so (during which the sensitivity to external pollution is poor, owing to hotter atmospheres; Koester et al. 2014; Barstow et al. 2014), a given white dwarf may exhibit metals at the observed frequency as a fraction of its remaining cooling lifetime (i.e. a duty cycle), or a subset of the entire population may stay constantly polluted, while others are never enriched by remnant planetary systems.
@@ 4.1 Lack of core dredge-up in DZ white dwarfs?
That being said, for a range metal abundances in DZ white dwarfs it has been demonstrated that atmospheric carbon bands, that might otherwise result from dredge-up, can be suppressed or masked (Blouin, 2022). Updated modeling shows that the C2 Swan band opacity decreases significantly with increasing metal abundance associated with external pollution, and thus, in principle, a star that appears spectrally of type DQ, may in principle transform into spectral type DZ. While certainly plausible for a range of parameters, roughly one quarter of the modeled DZQ spectra appear detectable with sufficient S/N. Such an empirical test should discover if this potential mechanism actually operates.
@@
It is well known that some DQ white dwarfs exhibit carbon features in the optical, while others do so only in the ultraviolet, and that this can be ascribed to differences in helium mass fractions rather than effective temperature (Weidemann & Koester, 1995; Blouin et al., 2023). In fact, as noted in Table 1, several of the DQ targets were found to have an optical C2 feature for the first time in their UVES spectra (0806–661, 1708–147, 1837–619), while others only exhibit carbon features in the ultraviolet (1917–077, 2147+280, 2317–173) despite the new, highly-sensitive observations. Therefore, it is possible or even likely that there is a continuous distribution of helium layer masses (Koester et al., 2020), where thinner layers will lead to optical carbon features, modestly thicker layers result in ultraviolet carbon features only, and if sufficiently thick, no carbon will be effectively dredged.
@@ 4.2 Lack of circumstellar material orbiting DQ stars
The dearth of detected circumstellar material orbiting DQ white dwarfs may appear – superficially – redundant with the lack of atmospheric pollution, as there is a strong connection established between these two hallmarks of evolved planetary systems. As a rule for white dwarfs, photospheric heavy elements require ongoing or recent accretion; that is, pollution implies closely-orbiting circumstellar material, where it is understood that the bulk of debris disks are simply undetectable owing to sensitivity limits (Rocchetto et al., 2015; Bonsor et al., 2017). But the converse is not necessarily true, and planetary matter can exist in a wide range of orbits and sizes that would not necessarily lead to the chemical enrichment of the host star. Little is known about the architecture of planetary systems orbiting polluted white dwarfs outside of a few tens of R⊙ (Xu et al., 2013; Farihi et al., 2014), but there are at least two examples of the aforementioned scenario: there is a giant planet candidate, transiting the non-polluted white dwarf WD 1856+534 (spectral type DC; Vanderburg et al. 2020), and a 2500 AU, planetary-mass companion orbiting WD 0806-661 (spectral type DQ, Table 1; Luhman et al. 2012).
@@ 4.3 Observed lack of close DQ+dM binaries
Because DC and DQ white dwarfs share similar stellar properties – with the likely exception of the atmospheric helium mass fraction – one would expect a comparable number of each in the SDSS catalog, adjusted for their relative occurrence rates as field stars. Within the 20 pc sample, there are 29 DC and 15 DQ white dwarfs (Hollands et al., 2018), and if their frequency in unresolved (including short-period) binaries were similar, one would expect close to 35 DQ spectral types within the composite binary catalog. Therefore, the lack of DQ stars in this catalog is striking. While identifying a weak C2 5165 Å feature may be difficult in low S/N spectra, and in the presence of flux from an M dwarf companion, many DQ white dwarfs exhibit deep and strong Swan bands that are unmistakable (Dufour et al., 2005).
@@
To demonstrate the plausibility of this atmospheric transformation, consider a K, pure helium atmosphere white dwarf, which will manifest as a DB spectral type, and where the fully mixed outer layer has mass M⊙(Koester, 2009; Koester et al., 2020). If this white dwarf accretes solar composition wind from a companion at a modest rate of M⊙ yr-1 (Pyrzas et al., 2012; Parsons et al., 2012; Ribeiro et al., 2013), hydrogen will come to dominate by number within a fraction of a Myr. With sufficient sensitivity, a hydrogen(-enriched) atmosphere white dwarf accreting stellar wind should appear as a DAZ spectral type, and there are several well-documented cases among post-common envelope binaries (Zuckerman et al., 2003; Debes, 2006; Tappert et al., 2011).
However, for the aforementioned, pure helium atmosphere case, once the convection zone deepens due to further cooling, any superficial hydrogen layer may be easily overwhelmed. By the time such a white dwarf has cooled to K, the size of the fully mixed layer should be M⊙, where this applies both to DQ white dwarfs and their helium atmosphere DC star counterparts (Koester, 2009; Koester et al., 2020). With convection zones of this depth, it would take several Gyr or longer for the atmosphere to become dominated by hydrogen that is captured from a companion stellar wind, with a strong dependence on the binary separation (i.e. wind capture rate). Therefore, it is expected that some DC, DZ, and DQ white dwarfs should be hosts to post-common envelope, main-sequence companions. Indeed, there are a handful of such systems known to host DC stars (Nebot Gómez-Morán et al., 2011; Parsons et al., 2013), and at least one candidate DZ accreting metals in a stellar wind from an M dwarf (Fajardo-Acosta et al., 2016). But there are none with a DQ white dwarf.
@@ 4.4 Apparently low DQ white dwarf masses
The left-hand panel of Figure 5 plots the mass distribution for the sample of 221 classical DQ white dwarfs discussed in Section 3.2 (Koester & Kepler, 2019). The masses of DQ stars have been studied recently by two independent teams, based on sources with reliable parallax data from Gaia, finding that their average mass is roughly 0.05 M⊙ lower than those found for DA and DB stars (Coutu et al., 2019; Bédard et al., 2022), and also lower than DZ and DC white dwarfs (Koester & Kepler, 2019). There have been suggestions that there may be model imperfections that prevent better agreement between DQ stars and other white dwarfs, but to date, no other studies have considered enhanced mass loss that may result from binary evolution.
It should be noted that two iconic DQ white dwarfs in the 20 pc sample have masses derived independently of models, and these values are also lower than the typical M⊙ found for spectral type DA and DB white dwarfs (Genest-Beaulieu & Bergeron, 2019). In fact, the dynamical mass of Procyon B (0.59 M⊙) is in modest tension with the overall age estimated for this well-studied binary system (Liebert et al., 2013; Bond et al., 2015). Adopting a comprehensive, recent, and empirical initial-to-final mass relation, the predicted mass of Procyon B should be closer to 0.66 M⊙ (Cummings et al., 2018). Similarly, for the planet-hosting binary GJ 86, the DQ white dwarf has a mass of 0.54 M⊙ (Zeng et al., 2021). And while it is likely this system is relatively old (either thick disk or somewhat younger; Fuhrmann et al. 2014), any recent initial-to-final mass relation will yield a progenitor that should not have evolved off the main sequence in a Hubble time. Therefore, it is possible that the masses of DQ stars are actually low compared to other white dwarfs, and not a result of model shortcomings, but as a consequence of their evolutionary history.
Among the 221 DQ white dwarfs plotted in Figure 5, there are 23 objects with derived masses M⊙, where 16 of these have , and thus excellent astrometric data. It is often the case that (apparently) over-luminous white dwarfs are two unresolved stars with comparable brightness, and therefore Gaia DR3 was searched in a 5 arcsec radius around the positions of these 23 candidate low-mass DQ stars; none were found to be resolved into two sources. However, two were found to have renormalized unit weight error (RUWE) suggestive of a possible or likely non-single star: SDSS J094058.34+384350.2 (RUWE ) and SDSS J125245.44+194311.2 (RUWE ). Thus, the inferred low masses of these 23 DQ white dwarfs may be partly accounted for as binary light sources, but the bulk currently lack indications of duplicity.
@@
Given that all DQ stars must have atmospheric carbon by definition, and thus carbon cores, it may seem oxymoronic to speculate on the possibility of low-mass members of this class. Owing to the finite age of the Galaxy, white dwarfs with M⊙ cannot yet have formed by single star evolution. Indeed the majority of low-mass white dwarfs are found to be single- or double-lined binaries (Marsh et al., 1995; Brown et al., 2011; Rebassa-Mansergas et al., 2011), whose evolution is thought to be terminated prior to helium ignition, and thus such stars would have helium cores as opposed to carbon cores.
However, it is possible to form a low-mass carbon core in stars with M⊙, where helium ignites non-degenerately (Bauer & Kupfer, 2021). Binary interactions would then be required to truncate stellar evolution and further core growth, in order to later be observed as a low-mass white dwarf with a carbon core. While speculative, this is a distinct possibility for some DQ white dwarfs, as their binary properties are only weakly constrained at present, and their molecular bands are relatively insensitive to radial velocity changes induced by companions, especially compared to white dwarfs with atomic spectral lines. The confirmation of any DQ white dwarf with M⊙ would be a strong indication of a binary origin.
The expectations for single star evolution are illustrated in the right-hand panel of Figure 5, where main-sequence progenitor masses have been calculated using the previously mentioned initial-to-final mass relation. Given the low masses of the DQ white dwarfs, it is thus not surprising that the predicted main-sequence predecessors are also of relatively low mass, but this important detail has so far escaped attention in previous studies (Coutu et al., 2019; Koester & Kepler, 2019; Bédard et al., 2022). The age of thin disk stars should be less than around 8 Gyr (Bensby et al., 2014; Kilic et al., 2017; Sharma et al., 2019), but Figure 5 indicates that most single star parents of DQ white dwarfs would have main-sequence lifetimes longer than this. For the DQ stars plotted in the Figure, K and M⊙, where white dwarf cooling would add an additional 1.2 Gyr to their inferred total ages. This picture would imply that DQ stars are dominated by thick disk and halo members, which, as shown below, is inconsistent with their kinematical and companion metallicity properties.